(1 point) Let X1, , X109, be IID with mean 16 and variance 9. By the CLT, P(X 109 > 16.09) can be approximated by 1-0(d) where Φ is the cdf of the standard normal distribution and (Round you answers to 2 decimal places)
Answer #1

here for 109 variables; expected std error =sqrt(9/109)=0.2873

therefore P(Xbar>16.09)=1-P(X<bar<16.09)=1-P(Z<(16.09-16)/0.2873)=1-P(Z<0.31)=1-0.62

therefore d =0.62